# MLRC-Bench: Can Language Agents Solve Machine Learning Research Challenges?

### v1

@@ 3.3 Objective Evaluation Metrics
Human Idea + MLAB (gpt-4o) 0.5 -1.0 -4.9 2.2 12.3 6.8 8.8 3.5
CoI-Agent Idea (o1) + MLAB (gpt-4o) 0.4 -1.0 -4.9 0.1 39.4 11.8 4.0 7.1
Metric: Relative Improvement to Human.
Quantitative performance comparisons across competitions can be tricky, as each task may differ significantly in its intrinsic difficulty, and the official baseline may be weaker or stronger. To address this, we use the “Relative Improvement to Human” as our main leaderboard metric that convert each raw performance score $s_{\text{agent}}$ into a normalized score $s^{{}^{\prime}}_{\text{agent}}$, using a linear transformation (Burns et al., 2024; Wijk et al., 2024). Therefore, the score of the baseline solution will be 0, and the top human solution in competition is set to 100. Formally, the normalization is computed as:
We repeat each agent for 8 trials, and report the best normalized performance. Normalized scores can exceed 100 if an agent’s solution surpasses the top human solution, although we have not observed normalized scores higher than 100. Solutions performing worse than the baseline will be assigned a negative normalized score.
3.4 Evaluation Protocol
Our evaluation protocol is meticulously designed to prevent AI agents from test set overfitting. Agents will submit their implementation in the form of an edited codebase, particularly within their proposed method in the methods/ directory. Specifically, in a single trial, an agent can iteratively modify the codebase multiple times. We store snapshots of the codebase immediately after each change. Whenever an execution occurs on the development set, we record the resulting metrics and the name of evaluated method for that snapshot. At the end of this iterative development phase, we pick the snapshot with the highest development performance (effectiveness). We then evaluate the method contained in that snapshot on the test set [^2] Concretely, we execute the command python main.py --method best_dev_method --phase test. for our final result. This approach strictly follows standard ML practice and ensures reproducible experimentation. Future work may explore more sophisticated multi-objective selection criteria that additionally weigh runtime (efficiency) or lines of code (simplicity) of implementations.
@@ 4 Experiments and Results
To evaluate the capability of LLM agents in solving ML research tasks, we conduct comprehensive experiments across different agent scaffoldings and language models. Each agent trial is conducted either on a single NVIDIA Quadro RTX 8000 GPU with 48GB of memory (for llm-merging, backdoor-trigger and rainfall-pred tasks) or a Tesla V100 GPU with 16GB memory (for all other tasks), determined by the size of the base model used in each task. Unless otherwise specified, we perform 8 trials per configuration and report the best attempt.
@@ 4.1 Agent Scaffolding Comparison
In addition to allow agents to directly propose and implement ideas, we investigate whether providing AI-generated or human-sourced ideas can enhance agent performance. Due to computational cost limits, we first evaluate GPT-4o (Hurst et al., 2024) under three scaffolding configurations:
  • MLAB: We apply the MLAB agent developed by Huang et al. (2024a). It is a ReAct-style (Yao et al., 2023) agent that alternates between thinking (such as reflection, research planning, fact-checking) and taking actions (file system operations or Python script execution) to implement methods.
  • CoI-Agent Idea + MLAB: We augment MLAB with ideas generated by Chain-of-Ideas (CoI) (Li et al., 2024a), an LLM-based agent that structures relevant literature into progressive chains to enhance ideation. We choose a strong model, OpenAI’s o1  (Jaech et al., 2024), as the backbone for the ideation agent to generate more creative ideas.
@@
For all tasks except Rainfall Prediction, we allow MLAB agents a maximum of 50 steps and 5 hours per trial, following the protocol established by Huang et al. (2024a). However, for Rainfall Prediction, we extend these limits to 100 steps and 10 hours, accounting for the increased baseline training time. The results in Table 3 show that incorporating additional ideas—whether generated by AI or proposed by humans—does not consistently lead to performance gains. This highlights the limited effectiveness and generalizability of AI-generated ideas, and emphasize the challenges agents face in implementing even human-proposed solutions effectively.
@@ 4.2 Model Comparison
Taking MLAB as our major scaffold, we evaluate five prominent LLMs: Claude 3.5 Sonnet v2 (Anthropic, 2024), gemini-exp-1206 (Pichai, 2024), Llama 3.1 405B Instruct (Dubey et al., 2024), o3-mini-high (OpenAI, 2025) and GPT-4o (2024-11-20) (Hurst et al., 2024). The results in Table 3 demonstrate varying success rates across models and tasks. Gemini-exp-1206 performs best among all models, reaching an average 9.3% relative improvement to human. Claude 3.5 Sonnet V2 performs the best on most tasks except failing drastically on machine unlearning. The agent’s proposed algorithm likely failed on machine unlearning because it treated the removal of unwanted data and the preservation of useful knowledge as separate steps, rather than jointly optimizing both goals to maintain model performance. We provide a more detailed case study of its failure in Section 4.6.
@@
Table 3 also reveals that agent solutions’ performance gains remain modest compared to human solutions in many cases, if not degrading baseline performance. There are, however, a few notable exceptions. For instance, MLAB (gpt-4o) achieves a score of 47.5 on the rainfall prediction task, likely because similar solutions (e.g., variants of U-Net (Ronneberger et al., 2015)) are readily available online. In the backdoor-trigger task, the baseline GCG method (Zou et al., 2023) performs poorly—essentially making random predictions—thereby lowering the bar for agents to surpass it with more meaningful solutions. This substantial gap highlights the current limitations of AI agents in generating novel, effective methods, underscoring the need for further advances to match or surpass human-led research efforts.
@@ 4.3 Inference-Time Scaling on ML Research Tasks
Increasing inference-time compute via repeated sampling (Chan et al., 2024; Chen et al., 2024a; Brown et al., 2024) has been shown to boost LLM performance on reasoning and coding tasks. Here we explore how LLM research agents scale with more inference-time compute on both the idea and solution spaces. We sample 4 ideas for each task from CoI-Agent (Li et al., 2024a) and repeat MLAB agent for 8 trials to implement each idea into code. Figure 2 plots pass@$k$ (Chen et al., 2021), i.e., the probability that at least one of $k$ trials converges to a successful implementation, defined as the agent closes at least 5% of the gap between baseline and top human participant scores (Relative Improvement to Human, Section 3.3).
@@
Figure 2: We measure Pass@$k$ as we scale the number of trials and ideas, running MLAB for eight trials per idea. The total inference-time computes are equivalent among these points: $k=4$ for one-idea line, $k=2$ for two-idea line, $k=1$ for four-idea line, and $k=4$ for the remaining lines. For results breakdown on each task, please refer to Figure 9 in Appendix B. Our results indicate that 1) providing high-quality ideas—especially human-generated ones—significantly boosts an agent’s success rate across multiple attempts, 2) while varying the balance between idea exploration and exploitation under a fixed budget yields similar outcomes due to diminishing returns from repeated trials.
@@
Our results show that providing high-quality ideas enhances an agent’s ability to generate meaningful solutions when given multiple attempts, and human ideas appear to be more effective than those produced by AI. Furthermore, under a fixed inference budget, we did not observe a significant difference between allocating resources to idea exploration versus exploitation. For example, there is no significant pass@k difference between using 4 ideas with 2 trials per idea, 2 ideas with 4 trials per idea, and 1 idea with 8 trials per idea. This phenomenon likely occurs because once a high-quality idea is identified, the performance gains from additional trials tend to plateau, resulting in diminishing returns despite further exploitation.
We hypothesize that performance-informed tree-search that navigates the vast space of possible solutions (Jiang et al., 2025; Koh et al., 2024) or allocating more computational resources (Chan et al., 2024) could offer more promising scaling properties, and we leave a detailed exploration of these exciting directions for future work.
Figure 3: Radar plots of objective and subjective evaluations for agent-generated solutions across seven research tasks. Each dimension is normalized on a 1–5 scale, where higher values indicate better performance. Objective metrics include effectiveness, efficiency, and simplicity, which are highlighted in bold. The rest are subjective metrics, assessed by prompting an LLM as a judge. Notably, more effective solutions identified by agents tend to be more complex and time-consuming (e.g., in backdoor trigger recovery task). Additionally, overlapping scores in subjective dimensions suggest that LLM-based evaluation struggles to distinguish the research capabilities of different models.
Figure 4: Correlation heatmap between objective (x-axis) and subjective (y-axis) metrics for agent-generated solutions across all tasks. In this setting, code is included when prompting the LLM to evaluate subjective dimensions. No strong correlation is observed, suggesting that LLM-judged subjective metrics may not reliably indicate real-world impact.
Figure 5: We track the percentages of changes of performance, runtime, and lines of code compared to baseline across iterative refinement of implementations within a trial of LLM-based MLAB agent on the development set. Performance improvement is the higher the better, while increased runtime and lines of code are the lower the better. For results breakdown on each task, please refer to Figure 10 and 11 in Appendix B.
4.4 Subjective Evaluation with LLM-as-a-Judge
Our benchmark also facilitates the investigation on whether LLM-as-a-judge style evaluations can reliably assess the quality of research ideas by comparing subjective (LLM-judged) and objective metrics. As shown in Figure 1, we first prompt an LLM to explain each implementation’s underlying idea. [^3] We instruct the MLAB agent to include detailed comments in the code to enable faithful, post-hoc explanations. We choose OpenAI’s o1 (Jaech et al., 2024) model here because of its superior reasoning and coding capabilities. We then apply the rubric from prior work (Baek et al., 2024) to have the o1 model assign a 1–5 Likert score on five dimensions: validity, clarity, rigorousness, generalizability, and innovativeness. These scores are all the higher the better. The prompts used for this evaluation are shown in Appendix C.
Furthermore, we examine how the presence of code influences the assessments through two settings. (1) Without Code, in which the LLM judges only access the task description and the proposed idea; and (2) With Code, in which the judges also see the code implementation. We then compute Spearman’s correlation (Spearman, 1904) for each pair of objective and subjective metrics, using data from all valid implementations that include test-phase scores.
Figure 3’s radar plots provide a holistic view of agent performance across both subjective and objective dimensions. The plots show that while agents occasionally produce effective solutions, they often struggle to balance other criteria such as efficiency and simplicity. For instance, on the backdoor-trigger task, Claude 3.5 Sonnet V2 scores well on effectiveness but poorly on efficiency and simplicity, suggesting that agent-generated solutions tend to be more complex and time-consuming. Notably, agents generally underperform compared to the baseline when evaluated using objective metrics. However, when subjective metrics are used—where LLMs serve as judges—they often receive more favorable ratings. This discrepancy highlights a risk of overly optimistic conclusions when relying solely on subjective evaluations.
Figures 4 and 8 (in Appendix B) illustrate the correlation heatmaps for both settings. [^4] We find that removing the code leads to similar correlation results and does not significantly affect the conclusion we make. The overall correlations remain weak. For example, there is a near-zero correlation (-0.06) between innovativeness and effectiveness, implying that an agent’s ability to generate novel ideas, as judged by an LLM, does not necessarily equate to success in practical tasks. Consequently, our finding indicates that LLM-based evaluations alone are not a reliable proxy for real-world research impact. While LLM agents can certainly assist in generating creative ideas, relying solely on LLM-based evaluations to gauge agents’ progress in improving machine learning research may lead to misinterpretations. This again highlights the importance of employing objective metrics to ensure that proposed solutions are not only novel but also effective.
4.5 Implementation Process Analysis
In Figure 6, we looked at how MLAB (gemini-exp-1206) agent progress through different stages when tackling tasks, categorizing each step into stages like exploring the problem, evaluating initial performance, generating new ideas, coding solutions, debugging errors, refining solutions, and final evaluations. We perform stage classification using gpt-4o with the prompt in Appendix F. We found that agents typically begin by spending a lot of time understanding the problem environment, then quickly move into writing and editing code. Interestingly, they don’t spend many steps brainstorming new ideas before jumping into implementation. This leads to frequent debugging, as rapid coding often introduces mistakes. Towards the end, instead of repeatedly refining their solutions, agents often rush to final testing and submission, suggesting there’s room to improve by encouraging more thoughtful planning and iterative improvement.
We additionally perform analysis on logs of agent traces and show the types, frequencies, and resolution patterns of coding errors, as well as agent behavior trends and capability levels, in Appendix D. We highlight two takeaways here. First, a large proportion of environmental errors were attributable to incorrect tool argument issues, which can be conceptualized as instances where the model “hallucinates” or misidentifies expected argument names. Notably, these errors accounted for approximately 11.5% of all environment steps. Second, a further examination of the overall error correction performance revealed that, across all evaluated tasks, the MLAB (gemini-exp-1206) agent resolved only 17.2% of the total errors encountered. This modest correction rate underscores the persistent difficulty of achieving robust error handling.
4.6 Case Study
We present two case studies below to illustrate failed solutions implemented by LLM agents. Please see Appendix E for the concrete code implemented by AI agents.
Machine Unlearning Challenge: The goal is to develop efficient algorithms that enable a model to “forget” specific training data, such that the resulting model closely resembles one that was never trained on that data in the first place. The MLAB agent (claude-3-5-sonnet-v2) proposed a Gradient Ascent Unlearning Algorithm, a two-phase approach combining gradient ascent for forgetting and fine-tuning for retaining knowledge. Specifically, the algorithm first performs gradient ascent on the forget set to maximize loss (achieving unlearning) and then fine-tunes the model on the retain set to restore the desired knowledge. While this approach sounds promising in theory, it scored significantly lower than the baseline. We hypothesize that by separating the gradient ascent on the forget set and the fine-tuning on the retain set into two distinct phases, the model may not effectively balance these two conflicting objectives. In contrast, a joint optimization approach—where both objectives are optimized at each gradient update—might better balance the processes of “forgetting” and “retaining” knowledge.
4.7 Cost-Effectiveness Analysis
In Figure 7, we analyze the agents’ success rates in a cost-controlled setting, motivated by recent work (Kapoor et al., 2024) emphasizing the importance of jointly optimizing both performance and cost in agent design. [^5] We exclude the gemini-exp-1206 model from this figure because it was experimental and its pricing was unavailable at the time of writing. Llama 3.1 405b Instruct [^6] We estimate the API cost for Llama models based on Amazon Bedrock service pricing. offers the most favorable trade-off, achieving higher success rates than GPT-4o and Claude 3.5 Sonnet at a significantly lower cost.
Figure 7: We perform a cost-effectiveness analysis of various setups. On the x-axis, we plot API cost, where lower is better, and on the y-axis, we show reletive improvement to human (Section 3.3), where higher is better. Among the settings evaluated, Llama 3.1 405B with the MLAB scaffolding emerges as a Pareto-optimal setting that balances cost and performance improvement.
5 Conclusion
@@
MLRC-Bench draws upon the rigor of conference competitions to provide a scalable, objective, and realistic benchmark for evaluating LLM agents in scientific discovery tasks. By isolating the core workflow of method proposal and implementation, MLRC-Bench highlights whether an agent can innovate, code, and refine solutions to meet objective performance criteria. Grounding its tasks in well-designed ML conference competitions ensures that each challenge addresses significant research topics, is carefully scoped with robust evaluation protocols, and includes code and data for objective, scalable, and reproducible assessment. Our benchmark results show that MLRC-Bench presents a significant challenge for state-of-the-art LLMs and agent scaffoldings. By featuring modular tasks, well-defined evaluation metrics, tamper-proof processes, and ongoing updates when new suitable competitions are available, MLRC-Bench can evolve alongside the rapid pace of ML research and continuous support the pursuit of AI-assisted or automated scientific discovery.

### v2

@@ 3.4 Evaluation Protocol
w/ Human Idea 0.5 -1.0 -4.9 2.2 12.3 6.8 8.8 3.5
w/ CoI-Agent Idea (o1) 0.4 -1.0 -4.9 0.1 39.4 11.8 4.0 7.1
@@ 4 Experiments and Results
To evaluate the capability of LLM agents in solving ML research tasks, we conduct comprehensive experiments across different agent scaffoldings and language models. Each agent trial is conducted either on a single NVIDIA Quadro RTX 8000 GPU with 48GB of memory (for llm-merging, backdoor-trigger and rainfall-pred tasks) or a Tesla V100 GPU with 16GB memory (for all other tasks), determined by the size of the base model used in each task. Unless otherwise specified, we perform 8 trials [^4] The number of trials is limited to 8 due to budget constraints on API usage. per configuration and report the best attempt.
@@ 4.1 Agent Scaffolding Comparison
In addition to allow agents to directly propose and implement ideas, we investigate whether providing AI-generated or human-sourced ideas can enhance agent performance. Due to computational cost limits, we follow the practice of MLE-Bench [7] to evaluate a commonly used model for agentic tasks, GPT-4o [22], under three scaffolding configurations:
  • MLAB: We chose a general-purpose MLAB [20] as our primary agent for evaluation.555 Another candidate agent framework, AIDE [24] is tailored for Kaggle-style tasks with single-file solutions and tabular outputs, making it less generalizable to our repository-level coding with diverse output formats (e.g., images, tensors). Its tree-search-based design with performance-based rewards is promising but requires careful adaptation, which we leave to future work. It is a ReAct-style [53] agent that alternates between thinking (such as reflection, research planning, fact-checking) and taking actions (file system operations or Python script execution) to implement methods.
  • CoI-Agent Idea + MLAB: We augment MLAB with ideas generated by Chain-of-Ideas (CoI) [31], an LLM-based agent that structures relevant literature into progressive chains to enhance ideation. We choose a strong model, OpenAI’s o1 [23], as the backbone for the ideation agent to generate more creative ideas.
@@
For all tasks, MLAB agents are allowed a maximum of 50 steps and 5 hours per trial, except for Rainfall Prediction, where the limits are extended to 100 steps and 10 hours due to the longer baseline training time. As shown in Table 2, incorporating additional ideas—whether generated by AI or proposed by humans—does not consistently improve performance. This underscores the challenges agents face not only in generating high-quality ideas but also in effectively implementing them, even when the ideas originate from humans.
@@ 4.2 Model Comparison
Taking MLAB as our major scaffold, we evaluate five prominent LLMs: Claude 3.5 Sonnet v2 [1], gemini-exp-1206 [42] [^6] Gemini-exp-1206 is renamed to Gemini 2.0 Pro afterwards., Llama 3.1 405B Instruct [15], o3-mini-high [41] and GPT-4o (2024-11-20) [22]. The results in Table 2 show varying success rates across models and tasks. Among all models, Gemini-exp-1206 performs the best overall, closing 9.3% of the gap between baseline and top human participant scores. Claude 3.5 Sonnet V2 performs the best on most tasks but fails significantly on machine unlearning. A case study of its final solution (Appendix C.1) suggests that the failure stems from treating the removal of unwanted data and the preservation of useful knowledge as separate objectives, rather than jointly optimizing both to maintain model performance. Furthermore, our cost-effectiveness evaluation in Appendix C.2 identifies Llama 3.1 405B with MLAB as a Pareto-optimal setup balancing API cost and performance.
@@
Table 2 also reveals that agent solutions’ performance gains remain modest compared to human solutions in many cases, if not degrading baseline performance. There are, however, a few notable exceptions. For instance, MLAB (gpt-4o) achieves a score of 47.5 on the rainfall prediction task, likely because similar solutions (e.g., variants of U-Net [43]) are readily available online. In the backdoor-trigger task, the baseline GCG method [56] performs poorly—essentially making random predictions—thereby lowering the bar for agents to surpass it with more meaningful solutions. This substantial gap highlights the current limitations of AI agents in generating novel, effective methods, underscoring the need for further advances to match or surpass human-led research efforts.
Figure 2: We track the percentages of changes of performance, runtime, and lines of code compared to baseline across iterative refinement of implementations within a trial of LLM-based MLAB agent on the development set. Performance improvement is the higher the better, while increased runtime and lines of code are the lower the better. These figures show the averaged metrics across all tasks. For results breakdown on each task, please refer to Figure 9 and 10 in Appendix D. The takeaway is that agents tend to over-refine their solutions over time, leading to increased complexity and runtime without proportional performance gains.
4.3 Solution Development Analysis
We further analyze the agent traces (gemini-exp-1206 with MLAB) in Appendix F to understand its behavioral patterns and limitations. Two key insights emerge. First, a significant portion of action errors—11.5% of all steps—stem from incorrect tool arguments, where the model either hallucinates or misidentifies expected parameter names. This highlights the need to improve the agent’s tool-usage capabilities. Second, the agent was able to fix only 17.2% of the errors encountered during code execution, revealing the challenge of self-debugging in complex, repository-level ML codebases [12].
Figure 3: Radar plots of objective and subjective evaluations for agent-generated solutions across seven research tasks. Each dimension is normalized on a 1–5 scale, where higher values indicate better performance. Objective metrics include Effectiveness, Efficiency, and Simplicity (Simp.), which are highlighted in bold. The rest are subjective metrics, assessed by prompting o1 as a judge. Notably, more effective solutions identified by agents tend to be more complex and time-consuming (e.g., in backdoor trigger recovery). Additionally, overlapping scores in subjective dimensions suggest that LLM-based evaluation struggles to distinguish the research capabilities of different models.
4.4 Subjective Evaluation with LLM-as-a-Judge
Our benchmark also facilitates the investigation on whether LLM-as-a-judge style evaluations can reliably assess the quality of research ideas by comparing subjective (here as judged by o1) and objective metrics. As shown in Figure 1, we first prompt an LLM to explain each implementation’s underlying idea. [^7] We prompt the MLAB agent to include detailed comments in the code to enable faithful explanations. We choose OpenAI’s o1 [23] model here because of its superior reasoning and coding capabilities. We then apply the rubric from prior work [2] to have the o1 model assign a 1–5 Likert score on five dimensions: validity, clarity, rigorousness, generalizability, and innovativeness. These scores are all the higher the better. The prompts used for this evaluation are shown in Appendix E. Furthermore, we examine how the presence of code influences the assessments through two settings. (1) Without Code, in which the LLM judges only access the task description and the proposed idea; and (2) With Code, in which the judges also see the code implementation. We then compute Spearman’s correlation [46] for each pair of objective and subjective metrics, using data from all valid implementations that include test-phase scores.
Figure 4: Correlation heatmap between objective (x-axis) and subjective (y-axis) metrics for agent-generated solutions across all tasks. Code is included when prompting the LLM to evaluate subjective dimensions. No strong correlation is observed, suggesting that LLM-judged subjective metrics may not reliably indicate empirical performance gains.
Figure 3’s radar plots provide a holistic view of agent performance across both subjective and objective dimensions. The plots show that while agents occasionally produce effective solutions, they often struggle to balance other criteria such as efficiency and simplicity. For instance, on the backdoor-trigger task, Claude 3.5 Sonnet V2 scores well on effectiveness but poorly on efficiency and simplicity, suggesting that agent-generated solutions tend to be more complex and time-consuming. Notably, agents generally underperform compared to the baseline when evaluated using objective metrics. However, when subjective metrics are used and judged by LLMs, they often receive more favorable ratings. This discrepancy highlights a risk of overly optimistic conclusions when relying solely on subjective evaluations.
Figures 4 and 7 (in Appendix D) illustrate the correlation heatmaps for both settings. [^8] We find that removing the code leads to similar correlation results and does not significantly affect the conclusion we make. The overall correlations remain weak. For example, there is a near-zero correlation (-0.06) between innovativeness and effectiveness, implying that an agent’s ability to generate novel ideas, as judged by an LLM, does not necessarily equate to success in practical tasks. Consequently, our finding indicates that LLM-based evaluations alone are not a reliable proxy for real-world research impact. While LLM agents can certainly assist in generating creative ideas, relying solely on LLM-based evaluations to gauge agents’ progress in improving machine learning research may lead to misinterpretations. This again highlights the importance of employing objective metrics to ensure that proposed solutions are not only novel but also effective.
@@ 4.5 Inference-Time Scaling on ML Research Tasks
Figure 5: We measure Pass@$k$ as we scale the number of trials and ideas, running MLAB for eight trials per idea. The total inference-time computes are equivalent among these points: $k=4$ for one-idea line, $k=2$ for two-idea line, $k=1$ for four-idea line, and $k=4$ for the remaining lines. For results breakdown on each task, please refer to Figure 8 in Appendix D. Our results indicate that 1) providing high-quality ideas—especially human-generated ones—significantly boosts an agent’s success rate across multiple attempts, 2) while varying the balance between idea exploration and exploitation under a fixed budget yields similar outcomes due to diminishing returns from repeated trials.
Increasing inference-time compute via repeated sampling [7, 8, 4] has been shown to boost LLM performance on reasoning and coding tasks. Here we explore how LLM research agents scale with more inference-time compute on both the idea and solution spaces. We sample 4 ideas for each task from CoI-Agent [31] and repeat MLAB agent for 8 trials to implement each idea into code.
@@
Figure 5 plots pass@$k$ [9], i.e., the probability that at least one of $k$ trials converges to a successful implementation, defined as the agent closes at least 5% of the gap between baseline and top human participant scores (Relative Improvement to Human, Section 3.3). Our results show that providing high-quality ideas enhances an agent’s ability to generate meaningful solutions when given multiple attempts, and human ideas appear to be more effective than those produced by AI. Furthermore, under a fixed inference budget, we did not observe a significant difference between allocating resources to idea exploration versus exploitation. For example, there is no significant pass@k difference between using 4 ideas with 2 trials per idea, 2 ideas with 4 trials per idea, and 1 idea with 8 trials per idea. This phenomenon likely occurs because once a high-quality idea is identified, the performance gains from additional trials tend to plateau, resulting in diminishing returns despite further exploitation. We hypothesize that performance-informed tree-search that navigates the vast space of possible solutions [24, 29] or allocating more computational resources [7] could offer more promising scaling properties.
5 Conclusion and Limitations
@@
MLRC-Bench draws upon the rigor of conference competitions to provide a scalable, objective, and realistic benchmark for evaluating LLM agents in proposing and implementing novel algorithms that advance research on impactful topics. Our benchmark features modular tasks, objective evaluation metrics, tamper-proof evaluations, and ongoing updates as new suitable competitions become available. Our results show that MLRC-Bench presents a significant challenge for state-of-the-art LLMs and agent scaffoldings. Our analysis highlights the misalignment between the LLM-judged innovation and their actual performance on cutting-edge ML research problems. MLRC-Bench will evolve alongside the rapid pace of ML research and continuously support the pursuit of AI-assisted or automated scientific discovery.
Our work has certain limitations. First, the benchmark currently covers only seven competitions, though we plan to expand it with new tasks as outlined in Section 3.1. Second, our evaluation focuses solely on a general-purpose MLAB agent for code implementation, as some other agent frameworks are currently unsuitable for our repository-level coding tasks (see Section 4.1). Nonetheless, recent advancements in LLM agents, including improved debugging tools [54], inference-time scaling techniques [49, 29], and reinforcement learning-based fine-tuning [55, 25, 10, 45], have the potential to achieve substantially better performance at a lower cost than human researchers.